BMK Corporation will need 215,000 Canadian dollars (C$) in 90 days to cover a payable position. Currently, a 90-day call option with an exercise price of $.75 and a premium of $.01 is available. Also, a 90-day put option with an exercise price of $.73 and a premium of $.01 is available. LKL plans to purchase options to hedge its payable position. Assuming that the spot rate in 90 days is $0.72, what is the net amount paid, assuming LKL wishes to minimize its cost?
Answer #1

As the firm needs to pay C $ in 90 days time, the firm is short on the underlying asset (which is C $ in this case). Hence, the firm will have to purchase the underlying asset and the same can be achieved by entering into a call option deal. Further, the option will be exercised only if the expected 90-day spot price is higher than the exercise price of the call option.

90-Day Expected Spot = $ 0.72,

Call Exercise Price = $ 0.75 and Call Premium = $ 0.01

Therefore, Net Call Price = Call Exercise Price + Call Premium = 0.75 + 0.01 = $ 0.76

As is observable, the call's net price is higher than the expected 90-day spot price and hence the firm will not exercise the call option. However, the firm has already purchased the call and hence has incurred the call's premium as an expense. Therefore, even if the option remains unexercised the net cost per C $ is (0.72 + 0.01) = $ 0.73 / C $

Net Amount Paid = 215000 x 0.73 = $ 156950